OpenStudy (jazzyluvsu123):

What is fraction version of 4.8 @satellite73 or @timaashorty

OpenStudy (anonymous):

improper fraction, or mixed number?

OpenStudy (anonymous):

\[.8=\frac{8}{10}=\frac{4}{5}\] so you can make it a mixed number in one step

OpenStudy (anonymous):

4 4/5

OpenStudy (jazzyluvsu123):

ok thanks

OpenStudy (anonymous):

it is correct

OpenStudy (jazzyluvsu123):

ok

OpenStudy (anonymous):

simplify \[\frac{ 48 }{ 10 }= 4 \frac{ 8 }{ 10 } = 4 \frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

@mathmale please simplify your answer

OpenStudy (mathmale):

The decimal quantity 4.8 can be re-written as \[\frac{ 4*10 }{ 10 }=\frac{ 48 }{ 10 }\] and that can be reduced. The result is an "improper fraction," not a mixed number.

OpenStudy (anonymous):

@mathmale, did she ask for an improper fraction. Teachers now a days want you to simplify noob